Solve each of the following applications. Give all answers to the nearest thousandth. Geometry. The length of a rectangle is 1 cm longer than its width. If the diagonal of the rectangle is 4 cm, what are the dimensions (the length and the width) of the rectangle?

Answer by checkley75(3666)

this is a right triangle with an hypotenuse of 4 & width of x & the length of (x+1).
using the pathagorean therum we get
x^2+(x+1)^2=4^2
x^2+x^2+2x+1=16
2x^2+2x+1-16=0
2x^2+2x-15=0
using the quadratic equation we get:
X=(-2+-SQRT[2^2-4*2*-15])/2*2
X=(-2+-SQRT[4+120])/4
X=(-2+-SQRT[124])/4
X=(-2+-11.14)/4
X=(-2+11.14)/4
X=9.14/4
X=2.285 ANSWER.
X=(-2-1114)/4
X=-13.14/4
X=-3.285 ANSWER.